OpenStudy (anonymous):

HELP Solve 2logx+log4=2 without a calculator. i know its so simple, but i haven't done these in a while please help.

OpenStudy (jhannybean):

Remember your log rules: \[\log(a)+ \log(b) = \log(ab)\]\[\log(x^2) = 2\log(x)\]

OpenStudy (anonymous):

which one would i use for my problem ? the first one correct ?

OpenStudy (jhannybean):

Both.

OpenStudy (jhannybean):

4 is a perfect square, meaning \(4 = 2^2\). That's a hint.

OpenStudy (anonymous):

so x would be? i don't get it /

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @Jhannybean Remember your log rules: \[\log(a)+ \log(b) = \log(ab)\]\[\log(x^2) = 2\log(x)\] \(\color{blue}{\text{End of Quote}}\) using the first rule, let's let \(a=x\) and \(b= 4\) Now lets rewrite it to fit the log rule: \(\log(a^2) +\log(b) =2\) Can you rewrite the left hand side of the equation for me following the log rules ?

OpenStudy (anonymous):

log(x^2) + log(4)=2

OpenStudy (jhannybean):

And so when youre adding two log functions together, the functions multiply, giving you \[\log(4x^2)=2\]

OpenStudy (jhannybean):

Understanding that so far

OpenStudy (anonymous):

okay so the answer would be x=5 correct?

OpenStudy (jhannybean):

I believe so. \[\log((2x)^2)= 2\]\[2\log(2x) = 2\]\[\log(2x) = 1\]\[10^{\log(2x)} =10^1\]\[2x=10\]\[x=5~~ \checkmark\]

OpenStudy (anonymous):

okay thank you ! ima post another question right now :)